find the slope and the y-intercept for the line that passes trough (-11,1) and (12,2)
(the answer need to be in integers or simplified fractions)

Answer by Alan3354(69443)

find the slope and the y-intercept for the line that passes trough (-11,1) and (12,2)
------------
slope m = diffy/diffx = 1/23
Use y = mx + b and either point to find b, the y-intercept
2 = (1/23)*12 + b
b = 34/23